# What to do if language models disagree? Black-box model ensembling for textual and visual question answering

### v1

@@ Abstract
A diverse range of large language models (LLMs), e.g., ChatGPT, and visual question answering (VQA) models, e.g., BLIP, have been developed for solving textual and visual question answering tasks. However, both LLMs and VQA models encounter challenges when applied to task-specific datasets. Fine-tuning these models is either difficult, as it requires access via APIs, rendering them as black-boxes, or costly due to the need of tuning a large number of parameters. To address this, we introduce InfoSel, a data-efficient and lightweight ensemble method that learns to dynamically pick the winner from existing black-box models for predictions on both textual and multimodal visual question answering tasks. Unlike traditional ensemble models, InfoSel does not rely on prediction probabilities or confidences, which typically are not available in black-box models. Experimental results on four datasets demonstrate that our approach achieves an absolute increase of up to +5.27% in the F1-score compared to standalone LLMs. Remarkably, this improvement is achieved by utilizing only 1K training instances and 110M model parameters for training task-specific ensemble models.
1 Introduction
@@
Large language models (LLMs) have demonstrated remarkable proficiency across a wide range of tasks, predominantly attributed to their ability to comprehend instructions and tap into vast repositories of high-quality data (Bubeck et al., 2023; Laskar et al., 2023). For example, ChatGPT finds extensive utilization in daily textual question answering (TQA) tasks, rendering substantial convenience to a myriad of users (Touvron et al., 2023).111https://chat.openai.com/ For visual question answering (VQA) tasks, VQA models have exhibited exceptional versatility, primarily due to their capability to comprehend both visual and textual context (Gong et al., 2023).
@@
However, recent work (Kocoń et al., 2023; Laskar et al., 2023) indicates that LLMs, such as ChatGPT, fall short of state-of-the-art performance on task-specific datasets. Similarly, VQA models (Li et al., 2022, 2021b; Bao et al., 2022) face challenges when applied to specialized datasets due to the idiosyncrasies in the content, format or structure of these datasets (Arora et al., 2018). Unfortunately, fine-tuning LLMs (e.g., LLaMA-2-70b-chat (Touvron et al., 2023)) on task-specific data requires a large number of GPU hours. Alternatively, training smaller, task-specific models from scratch requires large amount of labeled data in order to achieve comparable performance. Furthermore, fine-tuning LLMs through proprietary APIs with self-uploaded labeled training data not only requires LLM experts’ knowledge but is also expensive.222https://platform.openai.com/docs/guides/fine-tuning/ These fine-tuned models further remain black-box, with restricted access to details regarding architectural intricacies, model weights, training data, and even prediction confidences.
@@
In order to address these computational and accessibility challenges associated with fine-tuning, we introduce a scalable ensemble method called InfoSel (Informed Selection). InfoSel allows for training with just a few task-specific samples and is lightweight, with only a few million parameters. Unlike current LM-ensemble methods (e.g., MetaQA (Puerto et al., 2021), LLM-Blender (Jiang et al., 2023)) which depend on the confidence scores or log probabilities and thus can not be applied to black-box models like GPT3.5 text-davinci models, InfoSel does not rely on such information and offers black-box ensembling. Furthermore, our proposed ensemble method incorporates task-specific optimization, allowing it to be easily adapted to different datasets, considering variations of both the inputs and predicted answers from the ensembled LLMs (base models). This contrasts with traditional ensemble methods such as OLA (Woods et al., 1997) and PageRank (Brin & Page, 2012), which are not adapted to task-specific particularities (e.g., different features) of different datasets. Finally, our method efficiently deals with multimodal inputs. Concretely, our results exhibit superior performance on multimodal VQA inputs compared to state-of-the-art LLM-Blender (Jiang et al., 2023) ensemble method which is designed to work exclusively with text. Table 1 compares our method with alternatives.
OLA/ | ||||
FT | LLM-Blender | PageRank | InfoSel | |
Task-specific | ✓ | ✓ | ✗ | ✓ |
Data-efficient | ✗ | ✗ | ✓ | ✓ |
Lightweight | ✗ | ✗ | ✓ | ✓ |
Black-box | ✗ | ✗ | ✓ | ✓ |
Multimodal | ✗ | ✗ | ✓ | ✓ |
Ensemble | ✗ | ✓ | ✓ | ✓ |
@@
At its core, InfoSel (see Figure 1) trains a lightweight ensemble model to dynamically identify the most accurate base model (i.e., LLM or VQA model) for a given input, which we refer to as the winner. This is achieved by designing a meta-level classification task considering all the base models as labels for every input. We designed and implemented two ensemble architectures for textual and visual QA tasks. Our first proposed architecture, InfoSel-TT, uses textual transformer (TT, 110M parameters) (Devlin et al., 2019) as the backbone to generate textual representation of the question with the predicted answers by base models, and linearly mapping it to predict the F1-scores of these answers. Although InfoSel-TT is straightforward and effective, it cannot handle multimodal data. To address this, we propose a second architecture named InfoSel-MT, where we incorporate a multimodal transformer (MT, 115M parameters) (Li et al., 2019) to generate fused contextual representations of a multimodal input (image, question, and the predicted answers). These fused representations are used to train a dense layer for selecting the winner. The challenge with this approach is the lack of exposure of the base models to new (unseen) labels appearing in the task-specific datasets. To address this, we fine-tune TT and MT models (FT-TT and FT-MT) separately to learn these new labels. The predictions of these fine-tuned models are fused with the output from InfoSel using a second, separately trained InfoSel∗ ensemble model. We experiment with and without InfoSel∗, as this component is considered optional when InfoSel is already performing well.
@@
We select three LLMs (ChatGPT, LLaMA-2-70b-chat and GPT3.5 text-davinci-003) and three VQA models (ALBEF (Li et al., 2021a), BLIP (Li et al., 2022) and VLMo (Bao et al., 2022)). These models are used as ensemble base models to provide answers for textual and visual QA tasks respectively. To demonstrate the data efficiency of the proposed architectures, we train them on a subsample of training data from public benchmark datasets and test on the corresponding full test data, a setting referred to as “Mini-*”. Specifically, the Mini-SDv2 and Mini-NQ are sampled from SQuAD v2 (Rajpurkar et al., 2018) and NQ-Open (Kwiatkowski et al., 2019) respectively, for textual QA tasks. On the other hand, the Mini-GQA and Mini-Viz datasets are sampled from GQA (Hudson & Manning, 2019) and VizWiz (Gurari et al., 2018) VQA datasets (see details in Section 4.1). Experimental results showcase improvements of the performance up to +5.27% on Mini-SDv2 with InfoSel and +31.63% on Mini-Viz with InfoSel∗ when compared to the ensembled base models.
@@
In summary, our contributions are: (1) InfoSel, a novel lightweight and data-efficient approach to ensemble black-box models without relying on access to model architecture, weights or prediction confidences for optimizing on task-specific datasets; (2) Assessment of the performance on textual and multimodal visual QA tasks, demonstrating gains of up to +5.27% with InfoSel and up to +31.63% with InfoSel∗compared to ensembled base models on four benchmark datasets; (3) A detailed analysis of data efficiency, demonstrating that InfoSel surpasses the performance of the leading base models with as few as 10 training samples.333Code and data will be released upon acceptance.
@@ 2 Related Work
Domain Adaptation. These methods aim to improve the performance of a model on a task-specific domain by leveraging knowledge from other domains (Zhou et al., 2022). Methods such as fine-tuning (Yosinski et al., 2014), feature adaptation (Long et al., 2015) and data augmentation (Choi et al., 2019) aim to improve the performance of standalone models and thus typically require large amounts of labeled training data or access to the model architecture and weights. InfoSel addresses this challenge by employing a lightweight ensemble model trained on a small amount of labeled data, and restricted to access exclusively the predictions produced by black-box base models.
Ensemble Learning. Ensemble methods generate and combine multiple learners (ML models) to address a particular ML task (Sagi & Rokach, 2018). Classical ensembling approaches like boosting (Schapire, 2013) and bagging (Breiman, 1996) are designed to train and combine a large number of individual models and are thus computationally expensive. Snapshot ensemble method (Huang et al., 2017) uses several local minima from one single model for ensembling, which requires full access to model weights and architecture. Stacking (Wolpert, 1992) uses a meta-learner to integrate the probabilities of the predictions from base models for providing the final output. A recent LLM ensembling method proposed by Jiang et al. (2023) uses log probabilities generated by LLMs to train LLM-Blender models. Similarly, Puerto et al. (2021) introduce MetaQA to select the best answer from multiple experts which requires both the knowledge of confidence score and base models’ training data. Other methods train their base models to avoid dataset biases (Han et al., 2021), while Xu et al. (2019) aim to learn joint feature embeddings across different domains. However, these methods require at least one piece of knowledge that the black-box models can not provide, including base models’ confidence scores or even training data (not available for LLMs like ChatGPT). InfoSel is designed to not rely on any of the above elements.
Black-box Models Ensembling. The black-box ensembling can be achieved by classical dynamic classifier selection methods, most notably OLA (Woods et al., 1997), which learns to rank the best classifier dynamically by its overall local accuracy in the nearest region of the input. Alternatively, majority voting (Chan & van der Schaar, 2022) and PageRank (Brin & Page, 2012) weight the predictions by their internal agreements. Yet, these methods are not designed for task-specific optimization and do not consider the information about inputs and predicted answers from base models simultaneously. To address this, InfoSel proposes a transformer-based setup that utilizes all the information available in the black-box setting, but not more, to enhance task-specific performance.
@@ 3 InfoSel Ensemble Training
Figure 1 illustrates the proposed InfoSel and InfoSel∗ frameworks to ensemble LLMs for TQA tasks (left), and VQA models for VQA tasks (right). We differentiate TQA components using LLMs and VQA components using VQA models by denoting them with superscripts and respectively. Similarly, to distinguish between InfoSel, InfoSel∗ and FT models used in TQA and VQA tasks, we add suffixes “-TT” and “-MT” respectively. For example, the InfoSel-MT model in Figure 1, refers to the InfoSel for the VQA task.
@@ 3.1 InfoSel Training for TQA
Before training InfoSel, we first perform the data preparation (top of Figure 1) for both training and testing. Next, we train InfoSel and FT models, which is a necessary step before training the InfoSel∗.
Data Preparation. First, we randomly sample content-question pairs and the corresponding ground truth answers from various benchmark datasets (refer to Section 4.1). Next, we build prompts following specific prompt rules (refer to Section 4.1). Using these prompts instead of plain text improves the LLMs’ answer quality (Bach et al., 2022).
We select (=3) black-box LLMs to generate answers on the prompts. The answer generated by on is denoted as (). Thereby, LLMs provide candidate answers for prompts. We calculate the word-level -scores (Rajpurkar et al., 2018) of all the candidate answers respectively for . These -scores serve as target to optimize the ensemble model:
The input for the ensemble training consists of texts. Each text is formed by concatenating with each individual answer predicted by a base model , . More formally, the input is:
@@
InfoSel-TT. We use a textual BERT-base (Devlin et al., 2019) transformer , ( denote trainable model parameters) as the backbone of InfoSel-TT. To achieve faster convergence, we load the pre-trained weights of bert-base-uncased model.444https://huggingface.co/google-bert/bert-base-uncased The input vector is passed to to generate sentence representations for each value in respectively. Thus, the sentence representation of from is:
@@
FT-TT. A potential limitation of using InfoSel-TT model only is the lack of exposure of the base models to new (unseen) labels appearing in the task-specific datasets. To address this, we fine-tune a separate lightweight TT model directly on the TQA datasets to learn these new labels. Specifically, the training objective is to locate the start and end token position of the answer from the context . We provide the token positions of as the target label, such that the model is optimized to classify each token in two classes (start/end token). This fine-tuned textual transformer model is referred to as FT-TT ().555The training scheme is adapted from https://huggingface.co/learn/nlp-course/chapter7/7?fw=pt with the additional option to allow the model to return empty answers for unanswerable questions. We denote the answer predicted by on as .
InfoSel∗-TT. This model performs a further ensemble training of FT-TT and InfoSel-TT models with the same training scheme and labeled training data as InfoSel-TT. We anticipate that the thus trained InfoSel∗-TT model () on the output of InfoSel-TT and the label finetuned FT-TT, will improve the ability to handle labels unseen by base models. As a result, we expect an improvement in the overall task-specific performance. The winner model selected by belong to .
@@ 3.2 InfoSel Training for VQA
Data Preparation. Given image-question pairs from dev data of VQA benchmark datasets, we use (=3) pre-trained VQA models to predict answers as follows: . We denote the ground truth answer for image-question pair as . Target labels for ensemble training are given by the accuracy scores of the candidate answers evaluated on :
The concatenation of question () with each of the candidate answers () obtained from the base models and the corresponding image () serves the input to our ensemble model InfoSel-MT:
InfoSel-MT. A Multimodal Transformer (MT, ) (Li et al., 2021b) is employed as the backbone for InfoSel-MT. Specifically, we first generate visual features of using a pre-trained R-CNN model () (Anderson et al., 2018). is composed of a vector of the image region features and the detected tags ( i.e., object labels of the image) (Li et al., 2021b). The concatenated question-answer pair and is then passed together with to MT () to generate a fused contextual representation :
@@
FT-MT. Similar to FT-TT, FT-MT composed of a trainable MT and a Multilayer Perceptron (MLP) is fine-tuned with the same training data as InfoSel-MT. Differently, FT-MT solves a multi-label classification task by classifying the fused contextual representation of (instead of like InfoSel-MT) and to a predefined answer list (labels). This list contains frequent answers from the training data. As a result, a trained FT-MT model () can learn to predict the unseen (new) labels (answers) contained in the task-specific datasets, but not in the pre-training data of the base models. denotes the answer predicted by over . The training scheme is adapted from (Li et al., 2021b).
InfoSel∗-MT. Similar to InfoSel∗-TT, InfoSel∗-MT model ensembles the FT-MT and InfoSel-MT models using the same training scheme as in InfoSel-MT. The winner model selected by belong to .

### v2

@@ Abstract
A diverse range of large language models (LLMs), e.g., ChatGPT, and visual question answering (VQA) models, e.g., BLIP, have been developed for solving textual and visual question answering tasks. However, fine-tuning these models is either difficult, as it requires access via APIs, rendering them as black-boxes, or costly due to the need of tuning a large number of parameters. To address this, we introduce InfoSel, a data-efficient ensemble method that learns to dynamically pick the winner from existing black-box models for predictions on both textual and multimodal visual question answering tasks. Unlike traditional ensemble models, InfoSel does not rely on prediction probabilities or confidences, which typically are not available in black-box models. Experimental results on four datasets demonstrate that our approach achieves an absolute increase of up to +5.19% in the F1-score compared to standalone LLMs, up to +31.63% in accuracy compared to VQA models, using only 1K training instances. Besides, InfoSel surpasses baseline ensemble methods in data efficiency. We perform an in-depth analysis of the importance of fusing diverse multimodal information in ensembling.
keywords:
[…]
@@
Large language models (LLMs) have demonstrated remarkable proficiency across a wide range of tasks, predominantly attributed to their ability to comprehend instructions and tap into vast repositories of high-quality data [8, 23]. For example, ChatGPT finds extensive utilization in daily textual question answering (TQA) tasks, rendering substantial convenience to a myriad of users [29]. Furthermore, for visual question answering (VQA) tasks, VQA models have exhibited exceptional versatility, primarily due to their capability to comprehend both visual and textual context [16].
@@
However, recent work [21, 23] indicates that LLMs, such as ChatGPT, fall short of state-of-the-art performance on task-specific datasets such as question answering. Similarly, VQA models [24, 27, 5] face challenges when applied to specialized datasets due to the idiosyncrasies in the content, format or structure of these datasets [3]. Unfortunately, fine-tuning LLMs (e.g., LLaMA-2-70b-chat [37]) on task-specific data requires a large number of GPU hours. Alternatively, training smaller, task-specific models from scratch requires a large amount of labeled data to achieve comparable performance. Furthermore, fine-tuning LLMs through proprietary APIs with self-uploaded labeled training data not only requires LLM experts’ knowledge but is also expensive.111https://platform.openai.com/docs/guides/fine-tuning/ These fine-tuned models further remain black-box, with restricted access to details regarding architectural intricacies, model weights, training data, and even prediction confidences.
@@
In order to address these computational and accessibility challenges associated with fine-tuning, we introduce a scalable ensemble method called InfoSel (Informed Selection). InfoSel allows for training with just a few task-specific samples. Unlike current LM-ensemble methods (e.g., MetaQA [30]) which depend on the confidence scores and thus can not be applied to black-box models like GPT3.5 text-davinci models, InfoSel does not rely on such information and offers black-box ensembling. Furthermore, our proposed ensemble method incorporates task-specific optimization, allowing it to be easily adapted to different datasets, considering variations of both the inputs and predicted answers from the ensembled LLMs (base models). This contrasts with traditional ensemble methods such as OLA [39] and PageRank [7], which are not adapted to task-specific particularities (e.g., different features) of different datasets. Finally, our method efficiently deals with multimodal inputs. Concretely, our results exhibit superior performance on multimodal VQA inputs compared to state-of-the-art PairRanker [20] ensemble method which is designed to work exclusively with text. An in-depth exploration of the importance of utilizing diverse information as fused inputs to enhance the performance of ensemble training.
Fine-tuning [41] | Pair-Ranker [20] | PageRank [7] | OLA [39] | InfoSel (ours) | |
Task-specific | ✗ | ✗ | |||
Data-efficient | ✗ | ✗ | |||
Black-box | ✗ | ||||
Multimodal | ✗ | ✗ | |||
Ensemble | ✗ |
@@
At its core, InfoSel (see Figure 1) trains a small-size ensemble model to dynamically identify the most accurate base model (i.e., LLM or VQA model) for a given input, which we refer to as the winner. This is achieved by designing a meta-level classification task considering all the base models as labels for every input. We designed and implemented two ensemble architectures for textual and visual QA tasks. Our first proposed architecture, InfoSel-TT, uses textual transformer (TT, 110M parameters) [13] as the backbone to generate a textual representation of the question with the predicted answers by base models. Although InfoSel-TT is straightforward and effective, it cannot handle multimodal data. To address this, we propose a second architecture named InfoSel-MT, where we incorporate a multimodal transformer (MT, 115M parameters) [26] to generate fused contextual representations of a multimodal input (image, question, and the predicted answers). These fused representations are used to train a dense layer to select the winner model. The challenge with this approach is the lack of exposure of the base models to new (unseen) labels appearing in the task-specific datasets. To address this, we fine-tune TT and MT models (FT-TT and FT-MT) separately to learn these new labels. The predictions of these fine-tuned models are fused with the output from InfoSel using a second, separately trained InfoSel∗ ensemble model. We experiment with and without InfoSel∗, as this component is considered optional when InfoSel is already performing well.
@@
We select three LLMs (ChatGPT, LLaMA-2-70b-chat and GPT3.5 text-davinci-003) and three VQA models (ALBEF [25], BLIP [24] and VLMo [5]). These models are used as ensemble base models to provide answers for textual and visual QA tasks respectively. To demonstrate the data efficiency of the proposed architectures, we train them on a subsample of training data from public benchmark datasets and test on the corresponding full test data. Experimental results showcase improvements in the performance up to +5.19% on textual QA with InfoSel and +31.63% on VQA with InfoSel∗ when compared to the ensembled base models.
In summary, our contributions are:
@@
InfoSel, a novel data-efficient approach to ensemble black-box models without relying on access to model architecture, weights or prediction confidences for optimizing on task-specific datasets;
Assessment of the performance on textual and multimodal visual QA tasks, demonstrating gains of up to +5.19% with InfoSel and up to +31.63% with InfoSel∗compared to ensembled base models on four benchmark datasets (Section 5.1);
A detailed analysis of data efficiency shows that InfoSel is more data efficient than state-of-the-art baseline ensemble methods. InfoSel surpasses the performance of the leading base models with as few as 10 training samples on one benchmark dataset (Section 5.1);
An in-depth exploration of the importance of utilizing diverse information (image, question, answers) as fused inputs to enhance the performance of ensemble training (Section 5.2 and 5.3).222Code and data are available at https://github.com/Yuuxii/Black-box-QA-Ensemling.
@@ 2 Related Work
Domain Adaptation. These methods aim to improve the performance of a model on a task-specific domain by leveraging knowledge from other domains [42]. Methods such as fine-tuning [41], feature adaptation [28] and data augmentation [10] aim to improve the performance of standalone models and thus typically require large amounts of labeled training data or access to the model architecture and weights.
Ensemble Learning. Ensemble methods generate and combine multiple learners (ML models) to address a particular ML task [33]. Classical ensembling approaches like boosting [34] and bagging [6] are designed to train and combine a large number of individual models and are thus computationally expensive. Stacking [38] uses a meta-learner to integrate the probabilities of the predictions from base models for the final output. Puerto et al. [30] introduce MetaQA to select the best answer from multiple experts which requires both the knowledge of confidence score and base models’ training data. Other methods train their base models to avoid dataset biases [18], while Xu et al. [40] aim to learn joint feature embeddings across different domains. However, these methods require at least one piece of knowledge that the black-box models can not provide, including base models’ confidence scores or even training data (not available for ChatGPT).
Black-box Models Ensembling. Black-box ensembling can be achieved by classical dynamic classifier selection methods, most notably OLA [39], which learns to rank the best classifier dynamically by its overall local accuracy in the nearest region of the input. Alternatively, majority voting [9] and PageRank [7] weight the predictions by their internal agreements. Yet, these methods are not designed for task-specific optimization and do not consider the information about inputs and predicted answers from base models simultaneously. To address this, InfoSel proposes a transformer-based setup that utilizes all the information available in the black-box setting, but not more, to enhance task-specific performance.
@@ 3 InfoSel Ensemble Training
Figure 1 illustrates the proposed InfoSel and InfoSel∗ frameworks to ensemble LLMs for TQA tasks (left), and VQA models for VQA tasks (right). We differentiate TQA components using LLMs and VQA components using VQA models by denoting them with superscripts and respectively. Similarly, to distinguish models used in TQA and VQA tasks, we add suffixes “-TT” and “-MT” respectively. For example, the InfoSel-MT model in Figure 1, refers to the InfoSel for the VQA task.
[…]
@@ 3.1 InfoSel Training for TQA
Data Preparation. First, we randomly sample content-question pairs and the corresponding ground truth answers from various benchmark datasets (refer to Section 4.1). Next, we build prompts following specific prompt rules (refer to Section 4.1). Using these prompts instead of plain text improves the LLMs’ answer quality [4]. We select (=3) black-box LLMs to generate answers on the prompts. The answer generated by on is denoted as (). Thereby, LLMs provide candidate answers for prompts. We calculate the word-level -scores [31] of all the candidate answers respectively for . These -scores serve as target to optimize the ensemble model:
The input for the ensemble training consists of texts. Each text is a joint string of and an answer predicted by a base model , . More formally, the input is:
@@
InfoSel-TT. We use a textual BERT-base [13] transformer , ( denote trainable model parameters) as the backbone of InfoSel-TT. To achieve faster convergence, we load the pre-trained weights of bert-base-uncased model. The input vector is passed to to generate sentence representations for each value in respectively. Thus, the sentence representation of from is:
@@
FT-TT. Using only the InfoSel-TT model may limit performance due to the base models’ lack of exposure to new (unseen) labels in the task-specific datasets. To address this, we fine-tune a separate lightweight TT model directly on the TQA datasets to learn these new labels. Specifically, the training objective is to locate the start and end token position of the answer from the context . We provide the token positions of as the target label, such that the model is optimized to classify each token in two classes (start/end token). This fine-tuned textual transformer model is referred to as FT-TT ().333The training scheme is adapted from https://huggingface.co/learn/nlp-course/chapter7/7?fw=pt with the additional option to allow the model to return empty answers for unanswerable questions. We denote the answer predicted by on as .
InfoSel∗-TT. This model performs a further ensemble training of FT-TT and InfoSel-TT models with the same training scheme and labeled training data as InfoSel-TT. We anticipate that the thus trained InfoSel∗-TT model on the output of InfoSel-TT and the label finetuned FT-TT, will improve the ability to handle labels unseen by base models. As a result, we expect an improvement in the overall task-specific performance. The winner model selected by InfoSel∗-TT belong to .
@@ 3.2 InfoSel Training for VQA
Data Preparation. Given image-question pairs from dev data of VQA benchmark datasets, we use (=3) pre-trained VQA models to predict answers as follows: . We denote the ground truth answer for an image-question pair as . Target labels for ensemble training are given by the accuracy scores of the candidate answers evaluated on :
The joint string of a question () with each of the candidate answers () obtained from the base models, and the corresponding image () serve the input to our ensemble model InfoSel-MT:
InfoSel-MT. A Multimodal Transformer (MT, ) [27] is employed as the backbone for InfoSel-MT. Specifically, we first generate visual features of using a pre-trained R-CNN model () [1]. is composed of a vector of the image region features and the detected tags ( i.e., object labels of the image) [27]. The joint string of question-answer pair and is then passed together to MT () to generate a fused contextual representation :
@@
FT-MT. Similar to FT-TT, FT-MT composed of a trainable MT and a Multilayer Perceptron (MLP) is fine-tuned with the same training data as InfoSel-MT. Differently, FT-MT solves a multi-label classification task by classifying the fused contextual representation of (instead of like InfoSel-MT) and to a predefined answer list (labels). This list contains frequent answers from the training data. As a result, a trained FT-MT model () can learn to predict the unseen (new) labels (answers) contained in the task-specific datasets, but not in the pre-training data of the base models. denotes the answer predicted by over . The training scheme is adapted from [27].
InfoSel∗-MT. Similar to InfoSel∗-TT, InfoSel∗-MT model ensembles the FT-MT and InfoSel-MT models using the same training scheme as in InfoSel-MT. The winner model selected by InfoSel∗-MT belong to .